OpenStudy (anonymous):

If f(x) = 4x + 1 and g(x) = 3x + 2, what is f(g(x)) ? 12x + 9 12x + 7 12x + 5 12x + 3

OpenStudy (anonymous):

\[f(g(x)) = f(3x+2) = 4(3x+2)+1 = 12x+9 \]

OpenStudy (anonymous):

thanks